# Vulnerability Coordination Under the Cyber Resilience Act

### v2

@@ 4 The CRA and Open Source
Many open source software (OSS) projects and their foundations were critical about the CRA during its policy-making. Some notable OSS foundations even raised an argument that online distribution of OSS might be blocked in the EU due to the CRA (see [12], among many others during the CRA’s policy-making). While the argument might have merely been a part of an aggressive lobbying strategy, some have used a term “regulatory anger” to describe the situation during the political negotiations, concluding that the new regulatory obligations for OSS projects are modest at best and relatively easy to fulfill [16]. The reason for the only modest implications originates from the fact that the CRA essentially only applies to commercial OSS projects. This restriction is clarified in recitals 17, 18, and 19. These recitals clarify that natural or legal persons who merely contribute to OSS projects are not within the regulation’s scope. Nor is a mere distribution of OSS in the CRA’s scope. Instead, the keyword is monetization. To quote from recital 19, the CRA only applies to OSS projects that “are ultimately intended for commercial activities”, including “cases where manufacturers that integrate a component into their own products with digital elements either contribute to the development of that component in a regular manner or provide regular financial assistance to ensure the continuity of a software product”. This clarification essentially means that the main target is OSS foundations; the Linux Foundation would be the prime example. However, none of the keywords—including monetization, ultimate intentions, regular contributions, and financial assistance—are actually defined in Article 3. This lack of definitions makes it unclear how the regulators and legal systems will interpret the situation. Regardless, even with vague definitions and potential disputes, the actual regulatory mandates placed upon OSS foundations are only light.
@@
It can be mentioned that the initial legislative proposal for the CRA, identified as COM/2022/454 final, was more stringent about OSS. Thus, perhaps the lobbying by OSS stakeholders was also successful. Having said that, there are two important points many OSS lobbyists missed during the negotiations. The first is that the CRA’s Article 25 entails a development of voluntary security attestation programmes for OSS projects to ensure compliance with the essential cyber security requirements. Such programmes can be interpreted to boost the security of open source software [16]. The second point stems from Article 13(6), which mandates that a vendor who discovers a vulnerability from a third-party software component, including an OSS component, must report the vulnerability to the natural or legal persons responsible for the component’s development or maintenance. Also this reporting obligation can be seen to improve the security of open source software too. It is also directly related to supply chains.
@@ 5 The CRA and Supply Chains
A SBOM is defined in Article 3(39) to mean “a formal record containing details and supply chain relationships of components included in the software elements of a product with digital elements”. In practice, SBOMs are primarily about software dependencies, whether managed through a package manager or bundled into a software product directly. These are also under active research particularly in software engineering [3, 18]. As motivated by the CRA’s recital 77, SBOMs are seen not only important for regulatory purposes but they are also perceived as relevant for those who purchase and operate software products. However, the same recital notes that vendors are not obliged to make SBOMs public, although they must supply these to market surveillance authorities upon request according to Article 13(25). Confidentiality must be honored in such cases. But SBOMs are not the only new obligation related to supply chains.
Historically, there was a problem for many vulnerability discoverers to find appropriate contact details about organizations and their natural persons responsible [25]. The problem and related obstacles likely still persist today, as hinted by experiences with large-scale coordinated vulnerability disclosure and the poor adoption rate of new initiatives such as the so-called security.txt for the world wide web [6, 15, 27]. To this end, paragraph 6 in Annex I mandates vendors to also provide contact details for reporting vulnerabilities. This obligation is important also for regulators; as was noted in Section 2, they should coordinate vulnerability handling, including disclosure, also in case multiple parties are involved. Without knowing whom to contact, let alone which components have been integrated, such multi-party coordination is practically impossible. The same point applies to the earlier note about obligations placed upon vendors who may discover vulnerabilities in third-party components they use. The obligation to provide contact details is recapitulated in paragraph 2 of Annex II.
@@ 6 Conclusion
The paper examined, elaborated, and discussed the implications from the EU’s new CRA regulation upon vulnerability coordination. When keeping in mind the paper’s framing only to vulnerability coordination, the CRA’s new obligations cannot be argued to be substantial—in fact, it can be argued that many of these obligations have already long belonged to security toolboxes of responsible vendors prioritizing cyber and software security. Nevertheless, there are some new implications and associated research possibilities worth briefly pinpointing.
@@
Of the new regulatory obligations and points raised thereto, (1) further research is required about the concept of actively exploited vulnerabilities. Such research might also generally reveal whether the existing empirical vulnerability research has been biased in a sense that serious vulnerabilities affecting actual deployments have perhaps been downplayed; with some notable exceptions [4], the focus has often been on reported vulnerabilities affecting software products, components, and vendors, not on their actual deployments [19]. On a more theoretical side of research, (2) further contributions are required also regarding the other fundamental concepts involved, including not only vulnerabilities but also security incidents and the so-called near misses. Then, regarding supply chains, (3) further research is required also on SBOMs. Although technical research is well underway in this regard, including with respect to a question of how to generate SBOMs automatically [20], many questions remain about applicability, usability, and related aspects, including for regulatory purposes. Although perhaps only partially related to the paper’s explicit theme, (4) a relevant research question would also be whether, how, and how well vendors can comply with the requirement to provide security patches, preferably in an automated manner. When recalling the CRA’s scope, the question is non-trivial. While providing automated security updates for Internet of things devices sounds like a reasonable and well-justified goal, the situation may be quite different in industrial settings [24]. It also remains generally unclear whether the CRA applies to physical security, including patching of software operating in air-gapped and related deployments. As always, (5) further policy-oriented research is also needed. Among other things, a close eye is needed to monitor and evaluate the regulation’s future implementation, adoption, administration, and enforcement.
@@
Regarding the CRA’s future implementation, (6) a particularly relevant research topic involves examining the upcoming technical standards upon which the essential cyber security requirements are based. This topic is important already because complying with these standards provides a presumption of conformity according to the CRA’s Article 27(1). The topic can be extended to other (7) conformity assessment procedures involved, including the mandatory third-party audits required for critical products and the CRA’s relation to the older cyber security certification schemes in the EU. Among other things, a good topic to examine would be the accreditation of conformity assessment bodies. It may be that the CRA will also entail a new labor market demand for cyber security professionals specialized into security testing, auditing, and related domains. At the same time, as also implicitly remarked in the CRA’s Article 10, there is allegedly a cyber security skill gap in Europe and elsewhere [13]. As the gap may correlate with education [21, 29], (8) it may also be worthwhile to further study the CRA and its future implementation with education and curricula in mind. Regarding adoption, (9) it is relevant to know which conformity assessment procedures vendors will prefer in the future; according to Article 32(1), there are four distinct procedures for conformity assessments under the CRA.
@@
Although it is too early to speculate what the CRA’s broader implications may be, three general points can be still raised. First, the larger regulatory cyber security package in the EU has expectedly raised some criticism that fragmentation and complexity have increased [9, 23]. The CRA has contributed to this increase. The CRA also interacts with the EU’s other(cyber security laws. In particular, as the NIS2 directive also imposes reporting obligations for vendors, synchronization between the CRA and NIS2 has been perceived as crucial [7]. Second, a notable limitation of the CRA is that it offers no legal guards against criminal or civil liability from disclosing vulnerabilities. Such guards have often been seen important in the literature [2, 5]. Third, it remains to be seen whether the CRA actually improves the cyber security of products. Although all points require further research, including rigorous evaluations, particularly the last point is fundamental in terms of the CRA’s explicit goals. According to the Act’s first recital, cyber insecurity affects not only the EU’s economy but also its democracy as well as the safety and health of consumers. Already against this backdrop, it is important and relevant to evaluate in the future how well the CRA managed to remove or at least mitigate cyber insecurity in the EU.

### v3

@@ 4 The CRA’s Vulnerability Coordination Requirements for Open Source Projects
Many open source software (OSS) projects and their foundations were critical about the CRA during its policy-making. Some notable OSS foundations even raised an argument that online distribution of OSS might be blocked in the EU due to the CRA (see [13], among many others during the CRA’s policy-making). While the argument might have merely been a part of an aggressive lobbying strategy, some have used a term “regulatory anger” to describe the situation during the political negotiations, concluding that the new regulatory obligations for OSS projects are modest at best and relatively easy to fulfill [17]. The reason for the only modest implications originates from the fact that the CRA essentially only applies to commercial OSS projects. This restriction is clarified in recitals 17, 18, and 19. These recitals clarify that natural or legal persons who merely contribute to OSS projects are not within the regulation’s scope. Nor is a mere distribution of OSS in the CRA’s scope. Instead, the keyword is monetization.
To quote from recital 19, the CRA only applies to OSS projects that “are ultimately intended for commercial activities”, including “cases where manufacturers that integrate a component into their own products with digital elements either contribute to the development of that component in a regular manner or provide regular financial assistance to ensure the continuity of a software product”. This clarification essentially means that the main target is OSS foundations; the Linux Foundation would be the prime example. However, none of the keywords—including monetization, ultimate intentions, regular contributions, and financial assistance—are actually defined in Article 3. This lack of definitions makes it unclear how the regulators and legal systems will interpret the situation. Regardless, even with vague definitions and potential disputes, the actual regulatory mandates placed upon OSS foundations are only light.
@@
It can be mentioned that the initial legislative proposal for the CRA, identified as COM/2022/454 final, was more stringent about OSS. Thus, perhaps the lobbying by OSS stakeholders was also successful. Having said that, there are two important points that many OSS lobbyists missed during the negotiations. The first is that the CRA’s Article 25 entails a development of voluntary security attestation programmes for OSS projects to ensure compliance with the essential cyber security requirements. Such programmes can be interpreted to boost the security of open source software [17]. The second point stems from Article 13(6), which mandates that a vendor who discovers a vulnerability from a third-party software component, including an OSS component, must report the vulnerability to the natural or legal persons responsible for the component’s development or maintenance. Also this reporting obligation can be seen to improve the security of open source software too. It is also directly related to supply chains.
@@ 5 The CRA and Supply Chains
A SBOM is defined in Article 3(39) to mean “a formal record containing details and supply chain relationships of components included in the software elements of a product with digital elements”. In practice, SBOMs are primarily about software dependencies, whether managed through a package manager or bundled into a software product directly. These are also under active research [3, 20]. As motivated by the CRA’s recital 77, SBOMs are seen not only important for regulatory purposes but they are also perceived as relevant for those who purchase and operate software products. However, the same recital notes that vendors are not obliged to make SBOMs public, although they must supply these to market surveillance authorities upon request according to Article 13(25). Confidentiality must be honored in such cases.
Yet, SBOMs are not the only new obligation related to supply chains. Historically, there was a problem for many vulnerability discoverers to find appropriate contact details about organizations and their natural persons responsible [29]. The problem and related obstacles likely still persist today, as hinted by experiences with large-scale coordinated vulnerability disclosure and the poor adoption rate of new initiatives such as the so-called security.txt for the world wide web [6, 16, 33]. To this end, paragraph 6 in Annex I mandates vendors to also provide contact details for reporting vulnerabilities. This obligation is important also for regulators; as was noted in Section 2, they should coordinate vulnerability handling, including disclosure, also in case multiple parties are involved. Without knowing whom to contact, let alone which components have been integrated, such multi-party coordination is practically impossible. The same point applies to the earlier note about obligations placed upon vendors who may discover vulnerabilities in third-party components they use. The obligation to provide contact details is recapitulated in paragraph 2 of Annex II.
@@ 6 Conclusion
The paper examined, elaborated, and discussed the implications from the EU’s new CRA regulation upon vulnerability coordination. When keeping in mind the paper’s framing only to vulnerability coordination, the CRA’s new obligations cannot be argued to be substantial—in fact, it can be argued that many of these obligations have already long belonged to security toolboxes of responsible vendors prioritizing cyber and software security. Nevertheless, (1) among the new legal obligations are mandatory reporting of actively exploited vulnerabilities and severe cyber security incidents for which (2) strict reporting deadlines are also imposed. On the public administration side, (3) the CRA strengthens the EU-level administration due to ENISA’s involvement and the establishment of a common European reporting infrastructure, including a new European vulnerability database. In addition, (4) administrative complexity is also increased due to the involvement of new public sector bodies, including market surveillance authorities in particular. Furthermore, (5) the new mandatory reporting obligations apply also to OSS foundations with commercial ties, but also commercial vendors must report vulnerabilities to OSS projects whose software they use. Finally, (6) the CRA imposes new requirements also for supply chain management, some of which are related to vulnerability coordination; the notable examples are mandates to roll out SBOMs and provide contact details.
7 Further Work
Of the new regulatory obligations and points raised thereto, further research is required about the concept of actively exploited vulnerabilities. Such research might also generally reveal whether the existing empirical vulnerability research has been biased in a sense that serious vulnerabilities affecting actual deployments have perhaps been downplayed; with some notable exceptions [4], the focus has often been on reported vulnerabilities affecting software products, components, and vendors, not on their actual deployments [22]. Another related topic worth examining is the theoretical and practical relation between KEVs and the CRA’s notion of actively exploited vulnerabilities.
On a more theoretical side of research, further contributions are required also regarding the other fundamental concepts involved. Regarding the exploitation theme, many further concepts have been introduced in recent research, among them a concept of likely exploited vulnerabilities [19]. A conceptual enlargement applies also to security incidents. Although attempts have been made to better understand and theorize how different incident types are related to each other in terms of EU law [32], it still remains generally unclear what exactly separates a severe incident from a “conventional”, non-severe incident or a significant incident from a severe incident. These and other related interpretation difficulties can be seen also as a limitation of the current paper.
@@ 7 Further Work
Then, regarding supply chains, further research is required also on SBOMs. Although technical research is well underway in this regard, including with respect to a question of how to generate SBOMs automatically [23], many questions remain about applicability, usability, and related aspects, including for regulatory purposes. Although perhaps only partially related to the paper’s explicit theme, a relevant research question would also be whether, how, and how well vendors can comply with the requirement to provide security patches, preferably in an automated manner. When recalling the CRA’s scope, the question is non-trivial. While providing automated security updates for Internet of things devices sounds like a reasonable and well-justified goal, the situation may be quite different in industrial settings [28]. It also remains generally unclear whether the CRA applies to physical security, including patching of software operating in air-gapped and related deployments. As always, further policy-oriented research is also needed. Among other things, a close eye is needed to monitor and evaluate the regulation’s future implementation, adoption, administration, and enforcement.
@@
Regarding the CRA’s future implementation, a particularly relevant research topic involves examining the upcoming harmonized standards upon which the essential cyber security requirements are based. This topic is important already because complying with these standards provides a presumption of conformity according to the CRA’s Article 27(1). The topic can be extended to other conformity assessment procedures involved, including the mandatory third-party audits required for critical products and the CRA’s relation to the older cyber security certification schemes in the EU. Among other things, a good topic to examine would be the accreditation of conformity assessment bodies. It may be that the CRA will also entail a new labor market demand for cyber security professionals specialized into security testing, auditing, and related domains. At the same time, as also implicitly remarked in the CRA’s Article 10, there is allegedly a cyber security skill gap in Europe and elsewhere [14]. As the gap may correlate with education [24, 36], it may also be worthwhile to further study the CRA and its future implementation with education and curricula in mind. Regarding adoption, it is relevant to know which conformity assessment procedures vendors will prefer in the future; according to Article 32(1), there are four distinct procedures for conformity assessments under the CRA.
@@
Although it is too early to speculate what the CRA’s broader implications may be, three general points can be still raised. First, the larger regulatory cyber security package in the EU has expectedly raised some criticism that fragmentation and complexity have increased [9, 27]. The CRA has contributed to this increase. The CRA also interacts with the EU’s other legal cyber security acts. In particular, as the NIS2 directive also imposes reporting obligations for vendors, synchronization between the CRA and NIS2 has been perceived as crucial [7]. Second, a notable limitation of the CRA is that it offers no legal guards against criminal or civil liability from disclosing vulnerabilities. Such guards have often been seen important in the literature [2, 5]. Third, it remains to be seen whether the CRA actually improves the cyber security of products. Although all points require further research, including rigorous evaluations, particularly the last point is fundamental in terms of the CRA’s explicit goals. According to the Act’s first recital, cyber insecurity affects not only the EU’s economy but also its democracy as well as the safety and health of consumers. Already against this backdrop, it is important and relevant to evaluate in the future how well the CRA managed to remove or at least mitigate cyber insecurity in the EU.